OpenStudy (ccchristian1):

Equation A: y = 2x + 4 Equation B: y = 3x + 1 Which of the following is a step that can be used to find the solution to the set of equations? 2x + 4 = 3x + 1 2x = 3x + 4 2x + 4 = 3x 2x + 1 = 3x + 4

OpenStudy (ccchristian1):

plz help me

OpenStudy (ccchristian1):

anyone ?

OpenStudy (lalipapa123):

i;ll try

OpenStudy (ccchristian1):

thank you

OpenStudy (lalipapa123):

what do you think the answer is?

OpenStudy (lalipapa123):

its first

OpenStudy (anonymous):

Tried solving it and boom you will have the answer immediately (no kidding)

OpenStudy (lalipapa123):

if my answer helped you medal me

OpenStudy (ccchristian1):

first one ?

OpenStudy (lalipapa123):

yeah

OpenStudy (lalipapa123):

if my answer helped you plzz medal me!

OpenStudy (ccchristian1):

okay

OpenStudy (ccchristian1):

i have another question

OpenStudy (lalipapa123):

yes plzzz?!

OpenStudy (anonymous):

The answer wasn't A :o

OpenStudy (ccchristian1):

whats the answer guys I'm confused

OpenStudy (lalipapa123):

then say me what it is!@Mateaus

OpenStudy (lalipapa123):

@Mateaus

OpenStudy (ccchristian1):

whats the right answer ?

OpenStudy (lalipapa123):

the answer is [a]

OpenStudy (ccchristian1):

you sure?

OpenStudy (ccchristian1):

@Mateaus

OpenStudy (ccchristian1):

what do you think @Mateaus

OpenStudy (lalipapa123):

yes i am syre about that

OpenStudy (lalipapa123):

if you substitute the value of 'y' from equ.1 in equ.2 you will get the same!

OpenStudy (ccchristian1):

ok thank you

OpenStudy (ccchristian1):

i got another question

OpenStudy (ccchristian1):

y = 3x + 3 y = x − 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

OpenStudy (lalipapa123):

To solve using substitution, you first have to choose one equation to work with, then isolate one variable in that equation. using the second equation, x - y = 3, I can solve for y by subtracting x from both sides of the eqn. - y = 3 - x Now multiply both sides by (-1) to solve for y y = -3 + x Now I substitute my solution for y ( -3 + x) into the other equation in place of y. 3x + y = 1 3x + (-3 + x) = 1 Group like terms and solve for x 3x + - 3 + x = 1 4x + -3 = 1 4x = 4 x = 1 Now you can solve for y by subsituting x = 1 into the solution for y: y = -3 + x y = -3 + x y = -3 + 1 y = -2

OpenStudy (ccchristian1):

so how would you do it on this question

OpenStudy (ccchristian1):

sorry for the questions, I'm going to medal you right after thus question

OpenStudy (lalipapa123):

kk

OpenStudy (ccchristian1):

so how would you do the steps for this question

OpenStudy (ccchristian1):

@lalipapa123

OpenStudy (anonymous):

eh oh yeah the first one was A. I was just messing around to show work

OpenStudy (lalipapa123):

i had gave the answer above naaa!

OpenStudy (ccchristian1):

oo lol

OpenStudy (lalipapa123):

yes? do you mean you want the steps for the first one?

OpenStudy (ccchristian1):

wait so all those steps that you just said its talking about the question i asked you

OpenStudy (ccchristian1):

so those steps are the correct answer ?

OpenStudy (lalipapa123):

yes

OpenStudy (lalipapa123):

but the steps are for the second question

OpenStudy (ccchristian1):

alright sorry i got confused

OpenStudy (lalipapa123):

wait the steps given by me are for the second question!

OpenStudy (lalipapa123):

thanks @ccchristian1